# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

### v1

@@ 3.3.1 Li-CO2 battery electrolyte transport properties
Where the self-diffusion coefficients are determined from the mean-square-displacement of center of mass presented in Figure S1, calculated from the MD simulations as follows [95]:
@@ 3.3.2 Li+ solvation structure from large scale MD simulation
We present the average Radial Distribution Function () and Coordination Number () profiles for BF, TFSI-, and EMIM+ ions, as well as DMSO solvent, surrounding a single Li+ ion. Additionally, we analyze the and between molecules of the same type, represented by their center-of-mass atoms. Li+ ions preferentially coordinate with molecules containing highly electronegative atoms, i.e., atoms with a negative partial charge. The flat curves in Figure 3 (c) indicate a weak interaction between Li+ and EMIM+. Consequently, Li+ primarily coordinates with BF, TFSI-, and DMSO. According to Figure 3 (a), (b), and (d), Li+ ions interact strongly with BF, TFSI-, and DMSO. The most probable distances between Li+ and the constituent atoms are: BF (Li+-F 2.47 Å, Li+-B 3.22 Å), TFSI- (Li+-O 2.57 Å, Li+-F 2.85 Å), and DMSO (Li+-O 2.45 Å). The Li+ cations in the electrolyte are more solvated by BF and DMSO than TFSI-, as evidenced by the higher peaks in the and the corresponding coordination number curves of Li+-BF. BF and DMSO are closer to the Li+ ions than TFSI- ions, playing a crucial role in the dissociation of LiTFSI salt. However, TFSI- can also access the first coordination shell of Li+ due to strong interactions between its O and F atoms and Li+. The presence of both BF and TFSI- facilitates favorable coordination between Li+ and F, potentially leading to the formation of a Li-F solid electrolyte interphase (SEI) that protects the Li-metal anode. Interactions between molecules of the same type are weak due to electrostatic repulsion, resulting in well-dispersed molecules throughout the simulated electrolyte. The average first shell coordination number values for a tagged Li+ cation were determined using a threshold distance of 3 Å, based on the range of the first peaks. Key findings from the coordination number analysis include: Li+-Li+ (, ), Li+-BF (, ), Li+-DMSO (), and Li+-TFSI- (, , ).
@@
The high ionic conductivity and transference number (Table 2) of the simulated electrolyte is derived from several key factors revealed by the solvation structure and interaction analyses Figure 3. Li+ cations are preferentially solvated by BF and DMSO, promoting LiTFSI salt dissociation and increasing free charge carriers. Weak Li+-TFSI- interactions facilitate easier Li+ transport. The presence of BF and TFSI- creates Li+ conducting channels through favorable Li+-F coordination. Well-dispersed ionic species due to weak same-type molecular interactions enhance efficient ion transport. Optimal coordination numbers (e.g., for BF, for DMSO) allow efficient Li+ transport while maintaining sufficient solvation. Low Li+-Li+ coordination () minimizes clustering, promoting ion mobility. Multiple solvating species create a dynamic environment, facilitating rapid solvating molecule exchange and enhancing Li+ mobility. The preferential solvation of Li+ by BF and DMSO allows for more free movement compared to the larger TFSI- anions, resulting in a higher proportion of the ionic current carried by Li+ and thus increasing the Li+ transference number.
@@ 3.4 Cell performance from finite element analysis
The simulated voltage-capacity curves of the Li-CO2 battery cell during the charge-discharge phase is presented in Figure 4 (a). This curve was derived from finite element analysis (FEA) using the atomistic simulation results shown in Figure 2 and Table 2, along with other parameters detailed in Tables S1, S2, and S3. The FEA results were compared with the experimental data obtained from the set-up described in the Supporting Information (SI), at a current density of 1 mA/cm2. The data illustrated in Figure 4 (a) indicate a strong correlation between the simulated results and the experimental voltage. During discharge, the cell potential exhibited a rapid decline from 3.1 V to a plateau of approximately 2.65 V, followed by a gradual reduction to the cutoff voltage of 2.2 V. During charging, the cell potential increased rapidly from 3.7 V to a plateau of 4.2 V, followed by a gradual rise to the cutoff voltage of 4.5 V. Furthermore, the specific capacity, calculated on the basis of the catalyst mass, was about 6200 mAh/g (2.48 mAh/cm2).
@@
Figure 4 (b) illustrates the impact of applied current density on the simulated discharge curves of Li-CO2 battery cell. The discharge capacity exhibits a substantial reduction, decreasing from 81,570 mAh/g at a low current density of 0.1 mA/cm2 to 6,200 mAh/g at a higher current density of 1 mA/cm2. This observation aligns with prior experimental results, indicating that current density has a significant influence on cell capacity [98]. Furthermore, the discharge voltage plateau decreases as the current density increases (Figure 4 (b)). The capacity decrease at elevated discharge rates can be attributed to limitations in CO2 transport through the cathode, which becomes saturated with electrolyte and unable to sustain the electrochemical reaction. Consequently, as the discharge rate escalates, CO2 reduction is confined to a localized area near the cathode-current collector interface. Furthermore, the rapid reduction in porosity due to Li2CO3 and C deposition on the active area’s surface further impedes CO2 transport into the cell, preventing complete utilization of the electrode’s porosity. Figures 5, S3 and S4 illustrate and analyze these phenomena in detail.
Figure S2 illustrates the overpotential behavior of the studied Li-CO2 battery cell during discharge, revealing a transition from a positive overpotential to a negative overpotential. The initial positive overpotential can be attributed to kinetic and mass transport limitations associated with the CO2 reduction reaction (ORR) at the cathode. This phenomenon is consistent with observations in other battery systems, such as vanadium-manganese flow batteries, where kinetic barriers result in a positive overpotential during discharge [99]. As the discharge progresses, the overpotential becomes negative, indicating a shift in the reaction dynamics. This transition is due to morphological changes in the cathode surface or discharge products (Figure 5 (a1) and (a2)) that could create a more favorable electrochemical environment [22]. Furthermore, as the discharge continues, the distribution of reactants and products becomes more balanced, alleviating initial transport limitations. This phenomenon is related to the morphological changes observed in the cathode surface during discharge (Figure 5 (a1) and (a2)). We note that Li et al. reported that the formation of (lithium oxalate) particles on the cathode surface created a more favorable electrochemical environment by providing additional active sites, improving conductivity and increasing the effective surface area [100]. These changes resulted in higher discharge capacity and improved cycling stability, underscoring the complexity of the discharge process and the importance of optimizing materials and design to enhance overall battery cell performance.
The observed behavior where the overpotential in the Li-CO2 battery cell during discharge decreases to 0 V as current density increases from 0.1 to 1 mA/cm2 can be explained by changes in the morphology and formation mechanism of Li2CO3/C (Figure 5 (a1) and (a2)). It has been found for the Li-air battery that at lower current densities, the deposited species (Li2O2) tends to form larger, toroidal structures, which can lead to higher overpotentials due to less efficient electron transport [101, 102, 103]. Similarly to what was observed for Li-O2 battery, as the current density increases in Li-CO2 battery, the deposited species, Li2CO3 and carbon form a thin film (see Figure 5 (a1) and (a2)), which maintains closer contact with the surface of the conductive catalyst, thus reducing the overpotential (Figure S2) because Li2CO3 has low electrical conductivity (Figure 2). Figure 4 (c) shows that the capacity of the studied Li-CO2 battery cell decreases exponentially with increasing current density, as can be described by the following equation [102]:
@@
where: = 7500 mAh/g is the baseline capacity, = 110000 mAh/g is a scaling factor, = 0.157 is a characteristic current density. At low current density (i<<0.157 ), the capacity tends towards its maximum value of 117500 mAh/g. This is the sum of the baseline capacity () and the additional capacity represented by . As the current density increases, there is a rapid decrease in the additional capacity provided by , resulting in a decrease in the total capacity . The parameter determines the rate at which the capacity decreases with increasing current density. A smaller would mean a faster decrease in capacity for a given increase in current density, while a larger would indicate a slower decrease. As the current density increases further, the capacity approaches the baseline value of 7500 mAh/g. This suggests that at high current densities, the battery’s capacity is primarily limited to its baseline capacity, as the contribution from the exponential term becomes negligible. The exponential decrease in capacity with increasing current density highlights a trade-off between power output and energy capacity [104]. At higher current densities (i>>0.157 ), the battery can deliver more power, but at the cost of reduced capacity. This is a critical consideration in applications where both high-power and high-capacity are required. The connection between cell voltage () and current density () during the discharge of the studied Li-CO2 battery cell in Figure 4 (d) is represented by the formula [102, 105]:
@@
where V and V. The parameter represents the slope of the voltage change with the logarithm of the current density/(mAcm-2), while represents the voltage at a reference current density of 1 mA/cm2. This formula indicates that as the current density increases, the cell voltage decreases logarithmically, a common characteristic in electrochemical cells due to increased polarization and decreased voltage at higher current densities [102, 105].
@@
Figure 5 (a1) and (a2) illustrate how porosity () and volume fraction of Li2CO3/C deposited species () change over space and time with current densities of i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S3 presents a comparison of these parameters at i=0.1, i=0.25, i=0.5, and i=1 mA/cm2. For high current density (i=1 mA/cm2), porosity decreases gradually, primarily near the CO2 feed side of the cathode (X/L=1), from its initial value of to 0.48 at the end of discharge. This decrease is due to the formation of a thin film of Li2CO3/C, represented by the gradual increase of from 0 to 0.25 near the cathode feed side. This film formation leads to electrode clogging, impeding efficient CO2 diffusion into the cathode. For low current density (i=0.1 mA/cm2), Li2CO3/C forms larger toroidal structures [101, 102, 103], evidenced by the gradual increase of across a wider region of the cathode from 0 to 0.64. This results in a distributed porosity decrease from 0.73 to 0.09, initially allowing more CO2 transport into the cathode. However, as discharge progresses, the substantial porosity decrease hinders CO2 diffusion even at low current densities. Figure 5 (b1) and (b2) present the CO2 concentration profiles within the cell during discharge at i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S4 compares concentration profiles at various current densities. At i=1 mA/cm2, the CO2 concentration () decreases sharply from its initial value of 4.8 mol/m3 to 0 mol/m3 in the porous cathode, indicating strongly impeded CO2 diffusion. At i=0.1 mA/cm2, the decrease is more gradual. The impediment to CO2 transport, which is inversely proportional to the current density, results from the accumulation of Li2CO3 and carbon on the active cathode surface. This accumulation reduces pore availability for the electrolyte, increases the amount of insulating species, and obstructs pores, collectively restricting CO2 transport and limiting cell capacity (Figure 4 (c)). To address these limitations, an ideal cathode should feature: large surface area for reaction sites and product deposition, large pores to facilitate active species transport and prevent clogging, high porosity to accommodate discharge product growth, and high electrical conductivity and stability. Such a cathode design could mitigate the observed limitations and potentially enhance the overall performance of Li-CO2 batteries [106].
@@ 4 Conclusion
DFT/AIMD calculations reveal high electrical conductivity in the cathode catalyst, while MD simulations show favorable electrolyte properties. FEA results demonstrate an exponential decrease in discharge capacity with increasing current density, from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, primarily due to CO2 transport limitations and reduced cathode porosity. Li2CO3/C deposit morphology transitions from larger toroidal structures at lower current densities to thin films at higher current densities, significantly impacting performance. The evolution of the overpotential during discharge reflects changing reaction dynamics, while CO2 concentration variation plays a crucial role in determining battery performance.
Our multiscale approach bridges atomic-level phenomena and cell-level performance, offering unique insights into spatiotemporal evolution of critical battery parameters and reducing the need for extensive prototyping. Future research should focus on enhancing CO2 transport, mitigating deposit formation, and optimizing cathode designs with large surface areas, optimal pore sizes, and high porosity. This framework accelerates the development of high-performance Li-CO2 batteries and sustainable energy storage technologies by providing insights into phenomena difficult to observe experimentally in real-time.
@@
This work was conceived by A. T. N.. M. L., S. C. S., and N. K. D., carried out the calculations, including first-principles DFT, AIMD results, large-scale MD simulations, and Finite Element Analyses. I. P., A. N., L. A. C., and A. S. K. were responsible for the experimental measurements. All authors contributed to the discussion of the results and the final manuscript.

### v5

@@ 3.3.1 Li-CO2 battery electrolyte transport properties
Where the self-diffusion coefficients are determined from the mean-square-displacement of center of mass presented in Figure 3 (b), calculated from the MD simulations as follows [95]:
@@ 3.3.2 Li+ solvation structure from large scale MD simulation
We present the average Radial Distribution Function () and Coordination Number () profiles for BF, TFSI-, and EMIM+ ions, as well as DMSO solvent, surrounding a single Li+ ion. Additionally, we analyze the and between molecules of the same type, represented by their center-of-mass atoms. Li+ ions preferentially coordinate with molecules containing highly electronegative atoms, i.e., atoms with a negative partial charge. The flat curves in Figure 3 (e) indicate a weak interaction between Li+ and EMIM+. Consequently, Li+ primarily coordinates with BF, TFSI-, and DMSO. According to Figure 3 (c), (d), and (f), Li+ ions interact strongly with BF, TFSI-, and DMSO. The most probable distances between Li+ and the constituent atoms are: BF (Li+-F 2.47 Å, Li+-B 3.22 Å), TFSI- (Li+-O 2.57 Å, Li+-F 2.85 Å), and DMSO (Li+-O 2.45 Å). The Li+ cations in the electrolyte are more solvated by BF and DMSO than TFSI-, as evidenced by the higher peaks in the and the corresponding coordination number curves of Li+-BF. BF and DMSO are closer to the Li+ ions than TFSI- ions, playing a crucial role in the dissociation of LiTFSI salt. However, TFSI- can also access the first coordination shell of Li+ due to strong interactions between its O and F atoms and Li+. The presence of both BF and TFSI- facilitates favorable coordination between Li+ and F, potentially leading to the formation of a Li-F solid electrolyte interphase (SEI) that protects the Li-metal anode. Interactions between molecules of the same type are weak due to electrostatic repulsion, resulting in well-dispersed molecules throughout the simulated electrolyte. The average first shell coordination number values for a tagged Li+ cation were determined using a threshold distance of 3 Å, based on the range of the first peaks. Key findings from the coordination number analysis include: Li+-Li+ (, ), Li+-BF (, ), Li+-DMSO (), and Li+-TFSI- (, , ).
@@
The high ionic conductivity and transference number (Table 2) of the simulated electrolyte is derived from several key factors revealed by the solvation structure and interaction analyses (Figure 3). Li+ cations are preferentially solvated by BF and DMSO, promoting LiTFSI salt dissociation and increasing free charge carriers. Weak Li+-TFSI- interactions facilitate easier Li+ transport. The presence of BF and TFSI- creates Li+ conducting channels through favorable Li+-F coordination. Well-dispersed ionic species due to weak same-type molecular interactions enhance efficient ion transport. Optimal coordination numbers (e.g., for BF, for DMSO) allow efficient Li+ transport while maintaining sufficient solvation. Low Li+-Li+ coordination () minimizes clustering, promoting ion mobility. Multiple solvating species create a dynamic environment, facilitating rapid solvating molecule exchange and enhancing Li+ mobility. The preferential solvation of Li+ by BF and DMSO allows for more free movement compared to the larger TFSI- anions, resulting in a higher proportion of the ionic current carried by Li+ and thus increasing the Li+ transference number.
@@ 3.4 Cell performance from finite element analysis
The simulated voltage-capacity curves of the Li-CO2 battery cell during charge/discharge cycles are shown in Figure 4(a). These curves were generated through finite element analysis (FEA) using input parameters derived from atomistic simulations (Figure 2, Table 2) and additional material properties detailed in Tables S1–S3 (Supporting Information). The FEA results were validated against experimental data collected at a current density of 1 mAh/cm2, as described in the Supporting Information (SI). A strong correlation between simulated and experimental voltages is observed. During discharge, the voltage profile exhibits three distinct regions: (1) an initial sharp voltage drop due to kinetic overpotential () [98], (2) a linear decline attributed to resistive losses () across the growing Li2CO3/C film (average plateau potential: 2.65 V), (3) a "sudden death" characterized by rapid voltage collapse after 5000 mAh/g. The latter two phenomena, previously described for Li-O2 batteries as charge-transport limitations through Li2O2 films [99], exhibit analogous behavior in Li-CO2 systems. During charging, the cell maintains a stable potential of approximately 4.2 V until reaching 6000 mAh/g, followed by a steep voltage rise. The system achieves a specific capacity of 6200 mAh/g (2.48 mAh/cm2) normalized to the catalyst mass.
@@
Figure 4 (b) illustrates the impact of applied current density of 1 mA/cm2, 0.5 mA/cm2, 0.25 mA/cm2, and 0.1 mA/cm2 on the simulated discharge curves of Li-CO2 battery cell and Figure S1 shows the same curves for current densities of 0.2 mA/cm2, 0.3 mA/cm2, 0.4 mA/cm2, 0.6 mA/cm2, 0.7 mA/cm2, 0.8 mA/cm2 and 0.9 mA/cm2. The discharge capacity exhibits a substantial reduction, decreasing from 81,570 mAh/g at a low current density of 0.1 mA/cm2 to 6,200 mAh/g at the high current density of 1 mA/cm2. This observation aligns with prior experimental results, indicating that current density has a significant influence on cell capacity [100]. Furthermore, the discharge voltage plateau decreases as the current density increases (Figure 4 (b)). The capacity decrease at elevated discharge rates can be attributed to limitations in CO2 transport through the cathode, which becomes saturated with electrolyte and unable to sustain the electrochemical reaction. Consequently, as the discharge rate increases, CO2 reduction is confined to a localized area near the cathode-current collector interface. Furthermore, the rapid reduction in porosity due to Li2CO3 and C deposition on the active area’s surface further impedes CO2 transport into the cell, preventing complete utilization of the electrode’s porosity.
Figure 4 (c) demonstrates that the capacity of the studied Li-CO2 battery cell decreases exponentially with increasing current density, which can be described by the following equation [101]:
@@
where: = 7500 mAh/g is the baseline capacity, = 110000 mAh/g is a scaling factor, = 0.157 is a characteristic current density. At low current density (i 0.157 ), the capacity tends towards its maximum value of 117500 mAh/g. This is the sum of the baseline capacity () and the additional capacity represented by . As the current density increases, there is a rapid decrease in the additional capacity provided by , resulting in a decrease in the total capacity . The parameter determines the rate at which the capacity decreases with increasing current density. A smaller would mean a faster decrease in capacity for a given increase in current density, while a larger would indicate a slower decrease. As the current density increases further, the capacity approaches the baseline value of 7500 mAh/g. This suggests that at high current densities, the battery’s capacity is primarily limited to its baseline capacity, as the contribution from the exponential term becomes negligible. The exponential decrease in capacity with increasing current density highlights a trade-off between power output and energy capacity [102]. At higher current densities (i 0.157 ), the battery can deliver more power, but at the cost of reduced capacity. This is a critical consideration in applications where both high-power and high-capacity are required. The connection between cell voltage () and current density () during the discharge of the studied Li-CO2 battery cell in Figure 4 (d) is represented by the formula [101, 103]:
@@
where V and V. The parameter represents the slope of the voltage change with the logarithm of the current density/(mA/cm2), while represents the voltage at a reference current density of 1 mA/cm2. This formula indicates that as the current density increases, the cell voltage decreases logarithmically, a common characteristic in electrochemical cells due to increased polarization and decreased voltage at higher current densities [101, 103]. The performance of Li-CO2 batteries is governed by a complex interplay of structural and operational parameters, including the initial cathode porosity, CO2 diffusion in the battery electrolyte, initial CO2 concentration in the cathode, and the kinetics of the cathodic reaction. These factors collectively influence key performance metrics, such as capacity and voltage plateau presented at a current density of 0.4 mA/cm2 in Figure S2 and Figure 4 (e) and (f). Our analysis reveals a strong positive correlation between porosity () and capacity (), as demonstrated by the best linear fit equation:
This trend is attributed to the increased surface area and improved CO2 accessibility in higher porosity structures, which facilitate more efficient electrochemical reactions. Similarly, the CO2 diffusion coefficient () and initial CO2 concentration () exhibit significant positive correlations with capacity, as shown by the best linear fit equations:
These results highlight the critical role of CO2 transport and availability in determining battery capacity. Enhanced diffusion and higher CO2 concentrations improve the accessibility of CO2 to active sites, thereby promoting more efficient electrochemical reactions. This is consistent with recent findings by Chen et al., who demonstrated that a Li-O2/CO2 battery can achieve a full discharge capacity of 6628 mAh/g and a long life of 715 cycles, surpassing the performance of pure Li-O2 batteries [104]. However, the cathodic reaction rate constant () shows a logarithmic relationship with both capacity and voltage plateau, as evidenced by:
This logarithmic relationship underscores the significant impact of reaction kinetics on battery performance, emphasizing the need for efficient catalysts to optimize reaction rates. Furthermore, both CO2 diffusion and concentration demonstrate positive correlations with the voltage plateau, as shown by:
These findings indicate that improving CO2 transport and availability not only enhances capacity but also stabilizes the voltage plateau, which is critical for achieving stable and efficient battery operation. A recent study demonstrated a high-performance Li-CO2 battery using a Cu(II) solid redox mediator, achieving a higher discharge voltage of 2.8 V and a lower charge potential of 3.7 V [25]. Therefore, optimizing the cathode porosity, CO2 diffusion and concentration, while improving reaction kinetics through advanced catalysts, are essential strategies for enhancing the performance of Li-CO2 batteries. Future research should focus on developing innovative electrode materials, electrolyte and catalytic systems to leverage these relationships and advance the practical application of Li-CO2 batteries as a sustainable energy storage technology. For instance, a recent study reported a high-rate Li-CO2 battery enabled by a 2D medium-entropy catalyst, which operates at a high current density of 5000 mA/g and capacity of 5000 mAh/g for up to 125 cycles, far exceeding previously reported values [32].
@@
Figure 5 (a1) and (a2) illustrate how porosity () and volume fraction of Li2CO3/C deposited species () change over space and time with current densities of i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S3 presents a comparison of these parameters at i=0.1, i=0.25, i=0.5, and i=1 mA/cm2. For high current density (i=1 mA/cm2), porosity decreases gradually, primarily near the CO2 feed side of the cathode (X/L=1), from its initial value of to 0.48 at the end of discharge. This decrease is due to the formation of a thin film of Li2CO3/C, represented by the gradual increase of from 0 to 0.25 near the cathode feed side. This film formation leads to electrode clogging, impeding efficient CO2 diffusion into the cathode. For low current density (i=0.1 mA/cm2), Li2CO3/C forms larger structures [105], evidenced by the gradual increase of across a wider region of the cathode from 0 to 0.64. This results in a distributed porosity decrease from 0.73 to 0.09, initially allowing more CO2 transport into the cathode. However, as discharge progresses, the substantial porosity decrease hinders CO2 diffusion even at low current densities. Figure 5 (b1) and (b2) present the CO2 concentration profiles within the cell during discharge at i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S4 compares concentration profiles at various current densities. At i=1 mA/cm2, the CO2 concentration () decreases sharply from its initial value of 4.8 mol/m3 to 0 mol/m3 in the porous cathode, indicating strongly impeded CO2 diffusion. At i=0.1 mA/cm2, the decrease is more gradual. The impediment to CO2 transport, which is inversely proportional to the current density, results from the accumulation of Li2CO3 and C on the active cathode surface. This accumulation reduces pore availability for the electrolyte, increases the amount of insulating species, and obstructs pores, collectively restricting CO2 transport and limiting cell capacity (Figure 4 (c)). The nucleation and growth of Li2CO3 significantly impact performance, as described in Figure 2 (c) and the detailed FEA model in section 2.4. Current density plays a crucial role in this process. At high current densities (1 mA/cm2), rapid nucleation near the CO2 feed side forms a thin Li2CO3/C film (Figure 5 (a1), (b1)). In contrast, low current densities (0.1 mA/cm2) promote slower growth, resulting in larger, more distributed Li2CO3 structures and a gradual decrease in porosity (Figure 5 (a2), (b2)).
To address these limitations, an ideal cathode should feature: large surface area for reaction sites and product deposition, large pores to facilitate active species transport and prevent clogging, high porosity to accommodate discharge product growth, and high electrical conductivity and stability. Such a cathode design could mitigate the observed limitations and potentially enhance the overall performance of Li-CO2 batteries [106].
@@ 4 Conclusion
DFT/AIMD calculations reveal high electrical conductivity of the cathode catalyst, while MD simulations show favorable electrolyte properties. FEA results demonstrate an exponential decrease in discharge capacity with increasing current density, from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, primarily due to CO2 transport limitations and reduced cathode porosity. The Li2CO3/C deposit morphology transitions from larger structures at lower current densities to thin films at higher current densities, significantly impacting performance.
Our multiscale approach bridges atomic-level phenomena and cell-level performance, offering unique insights into spatiotemporal evolution of critical battery parameters and reducing the need for extensive prototyping. Future research should focus on enhancing CO2 transport, improving discharge product decomposition, and optimizing cathode designs with large surface areas, optimal pore sizes, and high porosity. This framework accelerates the development of high-performance Li-CO2 batteries and sustainable energy storage technologies by providing insights into phenomena difficult to observe experimentally in real-time.
@@
This work was conceived by A. T. N.. M. L., S. C. S., and N. K. D., carried out the calculations, including first-principles DFT, AIMD results, large-scale MD simulations, and Finite element analysis. I. P., A. N., L. A. C., and A. S. K. were responsible for the experimental measurements. All authors contributed to the discussion of the results and the final manuscript.